# Lossy Beyond Diagonal Reconfigurable Intelligent Surfaces: Modeling and Optimization

### v1

@@ III-E Complexity Analysis
In this subsection, we analyze the computational complexity of the proposed algorithms. The MM-ADMM algorithm (Algorithm 1) employs a double-loop iterative design. In the inner loop, each iteration updates $\bar{\bm{\phi}}$, $\bar{\mathbf{y}}$ and $\mathbf{z}$, requiring a total of $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations. In the outer loop, the update of $\bar{\mathbf{f}}_{t}$ per iteration demands $\mathcal{O}\{M^{4}\}$ operations. Therefore, the overall complexity of Algorithm 1 is $\mathcal{O}\{I_{\mathrm{out}}(M^{4}+I_{\mathrm{in}}\bar{M}^{3}M^{3})\}$, where $I_{\mathrm{out}}$ and $I_{\mathrm{in}}$ represent the number of iterations in the outer and inner loops, respectively. For comparison, the low-complexity algorithm employs a single-loop ADMM framework, resulting in an overall complexity of $\mathcal{O}\{I_{\mathrm{sig}}\bar{M}^{3}M^{3}\}$, where $I_{\mathrm{sig}}$ denotes the number of iterations. This single-loop structure offers a significant reduction in computational complexity compared to Algorithm 1, with a trade-off in performance, which will be illustrated in Section V.
@@
Remark 2: The overall complexity of the MM-ADMM algorithm is identical for both group- and forest-connected BD-RISs. This is because, in forest-connected BD-RIS, each iteration requires $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations to update $\bar{\bm{\phi}}$. For forest-connected BD-RIS, the low-complexity algorithm further reduces overall complexity: updating $\bar{\mathbf{y}}$ now requires $\mathcal{O}\{M^{3}\}$ operations instead of $\mathcal{O}\{\bar{M}^{3}M^{3}\}$, leading to a total complexity of $\mathcal{O}\{I_{\mathrm{sig}}M^{3}\}$.
@@ IV-A System Model and Problem Formulation
We then consider an $M$-element BD-RIS-aided MU-MISO system with an $N$-antenna base station (BS) and $K$ single-antenna users. At the BS, it is assumed that exact and instantaneous channel state information (CSI) is available. Let $\mathbf{s}\triangleq[s_{1},\ldots,s_{K}]^{T}\in\mathbb{C}^{K\times 1}$ represent the transmit symbol vector with $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}_{K}$, and $\mathbf{W}\triangleq[\mathbf{w}_{1},\ldots,\mathbf{w}_{K}]\in\mathbb{C}^{N\times K}$ denote the precoder matrix at the BS. Here, $s_{k}\in\mathbb{C}$ and $\mathbf{w}_{k}\in\mathbb{C}^{N\times 1}$ are the transmit symbol and precoding vector for user $k$, respectively, where $k\in\mathcal{K}=\{1,\ldots,K\}$. At each user, the received signal is given by
$y_{k}={}$ $(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})\mathbf{w}_{k}s_{k}$ (52)
$+(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})\sum_{p\in\mathcal{K},p\neq k}\mathbf{w}_{p}s_{p}+n_{k},\forall k\in\mathcal{K},$
where $\mathbf{h}_{RT,k}\in\mathbb{C}^{N\times 1}$, $\mathbf{h}_{RI,k}\in\mathbb{C}^{M\times 1}$, and $\mathbf{H}_{IT}\in\mathbb{C}^{M\times N}$ denote the channel matrices from BS to user $k$, from BD-RIS to user $k$, and from BS to BD-RIS, respectively, and $n_{k}\sim\mathcal{CN}(0,\sigma_{k}^{2})$ denotes the AWGN for user $k$, where $k\in\mathcal{K}$. Define $\tilde{\mathbf{h}}_{k}\triangleq(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})^{H},\forall k\in\mathcal{K}$. The signal-to-interference-plus-noise ratio (SINR) for each user can be written as
$\gamma_{k}=\frac{\left|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{k}\right|^{2}}{\sum_{p\in\mathcal{K},p\neq k}\left|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{p}\right|^{2}+\sigma_{k}^{2}},\forall k\in\mathcal{K}.$ (53)
We aim to jointly design the transmit precoder and the BD-RIS matrix to maximize the sum-rate of the MU-MISO system, taking into account the losses in BD-RIS and the transmit power constraint at the BS. Hence, the problem is formulated as [^8] The symmetric constraint of $\bm{\Phi}$ in (54b) ensures that the symmetry of $\bm{\Phi}$ is maintained during the alternating updates of $\bar{\bm{\phi}}$ and $\bar{\mathbf{y}}$ in subsection IV-D.
@@
where $P$ is the total transmit power at the BS. Problem (54) is difficult to tackle due to the ratio terms within the logarithm in the objective function and the non-convex constraints introduced by the BD-RIS. To address this, in the following subsections, we first reformulate the objective function (54a) using FP theory [30, 31], converting (54) into a multi-block optimization problem, and then optimize all variables iteratively.
@@ IV-B FP Reformulation
To deal with the complicated sum-rate objective function, we employ the FP technique proposed in [30, 31], which, by introducing auxiliary variables $\nu_{k}$ and $\tau_{k},\forall k\in\mathcal{K}$, transforms the original problem (54) into the following form:
@@
where $\bm{\nu}\triangleq[\nu_{1},\ldots,\nu_{K}]^{T}\in\mathbb{R}^{K}$ and $\bm{\tau}\triangleq[\tau_{1},\ldots,\tau_{K}]^{T}\in\mathbb{C}^{K}$. Compared to the original problem (54), the above formulation is more tractable, as the objective function with respect to each variable is simple and concave. To utilize this property, we employ a BCD framework to iteratively update each variable block until convergence. In the following subsections, we will show the details for updating each block in (55).
@@ IV-C Update of $\bm{\nu}$, $\bm{\tau}$ and $\mathbf{W}$
Given $\bm{\Phi}$ and $\mathbf{W}$, the updates of variables $\bm{\nu}$ and $\bm{\tau}$ follow a standard procedure as given in [31]:
$\nu_{k}$ $=\gamma_{k},\forall k\in\mathcal{K},$ (56)
$\tau_{k}$ $=\frac{\sqrt{1+\nu_{k}}\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{k}}{\sum_{p\in\mathcal{K}}|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{p}|^{2}+\sigma_{k}^{2}},\forall k\in\mathcal{K}.$ (57)
@@
$\mathbf{w}_{k}=\left(\sum_{p\in\mathcal{K}}|\tau_{p}|^{2}\tilde{\mathbf{h}}_{p}\tilde{\mathbf{h}}_{p}^{H}+\lambda\mathbf{I}_{N}\right)^{-1}\sqrt{1+\nu_{k}}\tau_{k}\tilde{\mathbf{h}}_{k},\forall k\in\mathcal{K},$ (58)
where $\lambda\geq 0$ is the Lagrange multiplier associated with the power constraint (54d) and can be determined via a bisection search; see detailed discussions in [8, 30].
@@ IV-D Update of $\bm{\Phi}$
where $\mathbf{v}_{p}\triangleq\mathbf{H}_{IT}\mathbf{w}_{p}$ and $\mathbf{x}_{k}\triangleq\mathbf{H}_{IT}\sum_{p\in\mathcal{K}}\mathbf{w}_{p}\mathbf{w}_{p}^{H}\mathbf{h}_{RT,k}$, $\forall p,k\in\mathcal{K}$. Let $\mathbf{v}_{p}=[\mathbf{v}_{p,1}^{T},\ldots,\mathbf{v}_{p,G}^{T}]^{T}$ and $\mathbf{h}_{RI,k}^{H}=[\mathbf{h}_{RI,k,1}^{H},\ldots,\mathbf{h}_{RI,k,G}^{H}]$, we have $\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{v}_{p}=\sum_{g\in\mathcal{G}}\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{v}_{p,g}$. Let $\mathbf{x}_{k}=[\mathbf{x}_{k,1}^{T},\ldots,\mathbf{x}_{k,G}^{T}]^{T}$, and we have $\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{x}_{k}=\sum_{g\in\mathcal{G}}\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{x}_{k,g}$. We then apply the vectorization property to transform $\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{v}_{p,g}$ and $\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{x}_{k,g},\forall g\in\mathcal{G}$, as
@@
The challenge of solving problem (63) is equivalent to that of solving problem (20), where the same ADMM methodology in Section III-B.2 is utilized to handle constraint (54c). The only minor difference lies in the subproblem of variable $\bar{\bm{\phi}}$, which is still an unconstrained convex optimization and admits a closed-form solution given by
@@
IV-E Extension to Forest-Connected BD-RIS
The proposed ADMM solution for updating $\bm{\Phi}$ is formulated using group-connected BD-RIS, which can be adapted to forest-connected BD-RIS by applying the same modifications described in Section III-D.
IV-F Complexity Analysis
In this subsection, we provide a complexity analysis for the BCD iterative algorithm. In each iteration, the update of ${\bm{\nu}}$ and ${\bm{\tau}}$ require $\mathcal{O}\{K(M^{2}+MN+KN)\}$ operations, and the update of $\mathbf{W}$ has a complexity of $\mathcal{O}\{KN^{2}+I_{\mathrm{bs}}N^{3}\}$, where $I_{\mathrm{bs}}$ is the number of iterations for the bisection search. For the update of $\bm{\Phi}$, the calculation of $\mathbf{Q}$ incurs a complexity of $\mathcal{O}\{K(MN+N^{2})+\max(K^{2}\bar{M}^{4},K\bar{M}M)+\bar{M}^{3}M^{3}\}$, and each iteration of updating $\bar{\bm{\phi}}$, $\bar{\mathbf{y}}$, and $\mathbf{z}$ involves $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations. Hence, the overall complexity of the BCD algorithm is $\mathcal{O}\{I(I_{\mathrm{bs}}N^{3}+KMN+\max(KM^{2},K^{2}\bar{M}^{4})+\max(K^{2}N,KN^{2})+I^{\prime}\bar{M}^{3}M^{3})\}$, where $I$ and $I^{\prime}$ denote the number of iterations for BCD and updating $\bm{\Phi}$, respectively. The aforementioned complexity remains unchanged when adapted for forest-connected BD-RIS.
@@ V Performance Evaluation
In this section, we present the simulation results to evaluate the impact of losses in the group- and forest-connected reconfigurable admittance networks on the performance of BD-RIS-aided SISO and MU-MISO systems. The channels are modeled as a integration of large-scale fading and small-scale fading. Regarding the large-scale fading, we use the distance-dependent pathloss model given by $PL_{s}=\zeta_{0}(d_{s}/d_{0})^{-\varepsilon_{s}}$, $\forall s\in\{\mathrm{RT,RI,IT}\}$. Here, $\zeta_{0}$ denotes the path loss at the reference distance $d_{0}$, $d_{s}$ represents the distance associated with each link ($\mathrm{RT,RI,IT}$: transmitter-to-receiver, RIS-to-receiver, and transmitter-to-RIS, respectively), and $\varepsilon_{s}$ corresponds to the path loss exponent for each link. For small-scale fading, the transmitter-to-receiver link is modeled as a Rayleigh fading channel, whereas the RIS-to-receiver and transmitter-to-RIS links are modeled as Rician fading channels. The Rician factors for these links are denoted by $\kappa_{i}$, $\forall i\in\{\mathrm{RI,IT}\}$. In the following simulations for both SISO and MU-MISO systems, we set $\zeta_{0}=-30$ dB at a reference distance $d_{0}=1$ m for all channels. The distances are given as $d_{\mathrm{RT}}=52$ m, $d_{\mathrm{RI}}=2.5$ m and $d_{\mathrm{IT}}=50$ m. The path loss exponents are set to $\varepsilon_{\mathrm{RT}}=3.8$, $\varepsilon_{\mathrm{RI}}=2.2$ and $\varepsilon_{\mathrm{IT}}=2.5$. The Rician factors for the Rician fading channels are fixed at $\kappa_{\mathrm{RI}}=\kappa_{\mathrm{IT}}=2$ dB. The noise power is assumed as $\sigma^{2}=-80$ dBm for SISO systems, and $\sigma^{2}_{k}=-80$ dBm, $\forall k\in\mathcal{K}$, for MU-MISO systems.
(a) MM-ADMM algorithm
@@
Figure 3: (a): Average rate versus $L_{1}$ for SISO systems ($R=3\ \Omega$, $M=32$, $\bar{M}=4$, $P=20$ dBm); (b): Sum-rate versus $L_{1}$ for MU-MISO systems ($R=3\ \Omega$, $N=K=4$, $M=32$, $\bar{M}=4$, $P=20$ dBm)
In Fig. 3, we first illustrate the selection of the value for $L_{1}$ by plotting the average rate and sum-rate performance against different practical $L_{1}$ values, using MM-ADMM algorithm for SISO systems and BCD algorithm for MU-MISO systems, respectively. A practical range of parasitic inductance values is considered, and the performance of various BD-RIS architectures is evaluated within this range. For tree-connected and forest-connected architectures, the tridiagonal form is adopted. The results demonstrate that the value of $L_{1}$ in the circuit model has a significant impact on performance. Specifically, $L_{1}=6$ nH is identified as a reasonable choice, providing generally good performance in both SISO and MU-MISO systems. This value is therefore used in all subsequent simulations. The performance differences among BD-RIS architectures will be discussed in the next two subsections (V-A and V-B). From an optimization perspective, $L_{1}$ influences the vertical shift of the circle defined by (9) for each tunable admittance component, thereby affecting their feasible admittance values during optimization. From a practical design perspective, a $6$ nH inductor has sufficiently low inductance to avoid introducing non-negligible parasitic resistance in the circuit [32].
V-A BD-RIS-aided SISO System
(a) MM-ADMM algorithm

### v2

@@ III-D Complexity Analysis
In this subsection, we analyze the computational complexity of the proposed algorithms. The $M$-ADMM algorithm (Algorithm 1) employs a double-loop iterative design. In the inner loop, each iteration updates $\bar{\bm{\phi}}$, $\bar{\mathbf{y}}$ and $\mathbf{z}$, requiring a total of $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations. In the outer loop, the update of $\bar{\mathbf{f}}_{t}$ per iteration demands $\mathcal{O}\{M^{4}\}$ operations. Therefore, the overall complexity of Algorithm 1 is $\mathcal{O}\{I_{\mathrm{out}}(M^{4}+I_{\mathrm{in}}\bar{M}^{3}M^{3})\}$, where $I_{\mathrm{out}}$ and $I_{\mathrm{in}}$ represent the number of iterations in the outer and inner loops, respectively. For comparison, the low-complexity algorithm employs a single-loop ADMM framework, resulting in an overall complexity of $\mathcal{O}\{I_{\mathrm{sig}}\bar{M}^{3}M^{3}\}$, where $I_{\mathrm{sig}}$ denotes the number of iterations.
@@
Remark 4: The overall complexity of the $M$-ADMM algorithm is identical for both group- and forest-connected BD-RISs. This is because, in forest-connected BD-RIS, each iteration requires $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations to update $\bar{\bm{\phi}}$. For forest-connected BD-RIS, the low-complexity algorithm further reduces overall complexity: updating $\bar{\mathbf{y}}$ now requires $\mathcal{O}\{M^{3}\}$ operations instead of $\mathcal{O}\{\bar{M}^{3}M^{3}\}$, leading to a total complexity of $\mathcal{O}\{I_{\mathrm{sig}}M^{3}\}$.
Remark 5: The main advantage of the low-complexity algorithm lies in its substantially reduced computational complexity compared to the $M$-ADMM algorithm. Due to the non-convexity of the objective function, the exact $M$-ADMM solution is obtained in a double-loop structure, where the outer loop is to deal with the non-convex objective, while the inner loop is to deal with complex constraints of lossy BD-RIS using the ADMM framework. In contrast, the low-complexity algorithm directly constructs a convex objective function and thus needs only a single loop ADMM. This reduces the complexity from $\mathcal{O}\{I_{\mathrm{out}}(M^{4}+I_{\mathrm{in}}\bar{M}^{3}M^{3})\}$ to $\mathcal{O}\{I_{\mathrm{sig}}\bar{M}^{3}M^{3}\}$ for group-connected BD-RIS, and further to $\mathcal{O}\{I_{\mathrm{sig}}M^{3}\}$ for forest-connected BD-RIS. However, since it solves the approximate problem (40) rather than the original power maximization problem, performance degradation may occur when the approximation is inaccurate, e.g., for D-RIS. This complexity–performance trade-off will be quantified in Section V.
@@ IV-A System Model and Problem Formulation
We then consider an $M$-element BD-RIS-aided MU-MISO system with an NN-antenna base station (BS) and $K$ single-antenna users. At the BS, it is assumed that exact and instantaneous channel state information (CSI) is available. Let $\mathbf{s}\triangleq[s_{1},\ldots,s_{K}]^{T}\in\mathbb{C}^{K\times 1}$ represent the transmit symbol vector with $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}_{K}$, and $\mathbf{W}\triangleq[\mathbf{w}_{1},\ldots,\mathbf{w}_{K}]\in\mathbb{C}^{N\times K}$ denote the precoder matrix at the BS. Here, $s_{k}\in\mathbb{C}$ and $\mathbf{w}_{k}\in\mathbb{C}^{N\times 1}$ are the transmit symbol and precoding vector for user $k$, respectively, where $k\in\mathcal{K}=\{1,\ldots,K\}$. At each user, the received signal is given by
$y_{k}=$ $~(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})\mathbf{w}_{k}s_{k}$ (49)
$+(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})\sum_{p\in\mathcal{K},p\neq k}\mathbf{w}_{p}s_{p}+n_{k},~\forall k\in\mathcal{K},$
where $\mathbf{h}_{RT,k}\in\mathbb{C}^{N\times 1}$, $\mathbf{h}_{RI,k}\in\mathbb{C}^{M\times 1}$, and $\mathbf{H}_{IT}\in\mathbb{C}^{M\times N}$ denote the channel matrices from BS to user $k$, from BD-RIS to user $k$, and from BS to BD-RIS, respectively, and $n_{k}\sim\mathcal{CN}(0,\sigma_{k}^{2})$ denotes the AWGN for user $k$, where $k\in\mathcal{K}$. Define $\tilde{\mathbf{h}}_{k}\triangleq(\mathbf{h}_{RT,k}^{H}+\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{H}_{IT})^{H},~\forall k\in\mathcal{K}$. The signal-to-interference-plus-noise ratio (SINR) for each user can be written as
$\gamma_{k}=\frac{\left|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{k}\right|^{2}}{\sum_{p\in\mathcal{K},p\neq k}\left|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{p}\right|^{2}+\sigma_{k}^{2}},~\forall k\in\mathcal{K}.$ (50)
We aim to jointly design the transmit precoder and the BD-RIS matrix to maximize the sum-rate of the MU-MISO system, taking into account the losses in BD-RIS and the transmit power constraint at the BS. Hence, the problem is formulated as [^9] The symmetric constraint of $\bm{\Phi}$ in (51b) ensures that the symmetry of $\bm{\Phi}$ is maintained during the alternating updates of $\bar{\bm{\phi}}$ and $\bar{\mathbf{y}}$ in Section IV-D.
@@
where $P$ is the total transmit power at the BS. Problem (51) involves the same lossy BD-RIS constraints as problem (12) in the SISO case. In contrast to the SISO systems, where only the BD-RIS matrix is optimized, we should jointly design the transmit precoder and BD-RIS matrix in MU-MISO systems. Furthermore, the ratio terms within the logarithm of the objective function in problem (51) complicates the optimization. To tackle these challenges, in the following subsections, we first reformulate the objective function (51a) using FP theory [FPI, FPII], converting (51) into a multi-block optimization problem, and then optimize all variables iteratively.
@@ IV-B FP Reformulation
To deal with the complicated sum-rate objective function, we employ the FP technique proposed in [FPI, FPII], which, by introducing auxiliary variables $\nu_{k}$ and $\tau_{k},~\forall k\in\mathcal{K}$, transforms the original problem (51) into the following form:
@@
where $\bm{\nu}\triangleq[\nu_{1},\ldots,\nu_{K}]^{T}\in\mathbb{R}^{K\times 1}$ and $\bm{\tau}\triangleq[\tau_{1},\ldots,\tau_{K}]^{T}\in\mathbb{C}^{K\times 1}$. Compared to the original problem (51), the above formulation is more tractable, as the objective function with respect to each variable is simple and concave. To utilize this property, we employ a BCD framework [^10] The main difference between our BCD algorithm and the one in [weighted_sum_rate_ris_Guo] lies in the update of $\bm{\Phi}$. The considered lossy BD-RIS must satisfy the complex constraints in (17b) – (17d), rather than the simple unit-modulus constraint of D-RIS in [weighted_sum_rate_ris_Guo]. Hence, the algorithm in [weighted_sum_rate_ris_Guo] is not applicable here. to iteratively update each variable block until convergence. In the following subsections, we will show the details for updating each block in (52).
@@ IV-C Update of $\bm{\nu}$, $\bm{\tau}$ and $\mathbf{W}$
Given $\bm{\Phi}$ and $\mathbf{W}$, the updates of variables $\bm{\nu}$ and $\bm{\tau}$ follow a standard procedure as given in [FPII]:
$\nu_{k}$ $=\gamma_{k},~\forall k\in\mathcal{K},$ (53)
$\tau_{k}$ $=\frac{\sqrt{1+\nu_{k}}\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{k}}{\sum_{p\in\mathcal{K}}|\tilde{\mathbf{h}}_{k}^{H}\mathbf{w}_{p}|^{2}+\sigma_{k}^{2}},~\forall k\in\mathcal{K}.$ (54)
@@
$\mathbf{w}_{k}=\left(\sum_{p\in\mathcal{K}}|\tau_{p}|^{2}\tilde{\mathbf{h}}_{p}\tilde{\mathbf{h}}_{p}^{H}+\lambda\mathbf{I}_{N}\right)^{-1}\sqrt{1+\nu_{k}}\tau_{k}\tilde{\mathbf{h}}_{k},~\forall k\in\mathcal{K},$ (55)
where $\lambda\geq 0$ is the Lagrange multiplier associated with the power constraint (51d) and can be determined via a bisection search; see detailed discussions in [hybrid_hongyu, FPI].
@@ IV-D Update of $\bm{\Phi}$
where $\mathbf{v}_{p}\triangleq\mathbf{H}_{IT}\mathbf{w}_{p}$ and $\mathbf{x}_{k}\triangleq\mathbf{H}_{IT}\sum_{p\in\mathcal{K}}\mathbf{w}_{p}\mathbf{w}_{p}^{H}\mathbf{h}_{RT,k}$, $\forall p,k\in\mathcal{K}$. Let $\mathbf{v}_{p}=[\mathbf{v}_{p,1}^{T},\ldots,\mathbf{v}_{p,G}^{T}]^{T}$ and $\mathbf{h}_{RI,k}^{H}=[\mathbf{h}_{RI,k,1}^{H},\ldots,\mathbf{h}_{RI,k,G}^{H}]$, we have $\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{v}_{p}=\sum_{g\in\mathcal{G}}\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{v}_{p,g}$. Let $\mathbf{x}_{k}=[\mathbf{x}_{k,1}^{T},\ldots,\mathbf{x}_{k,G}^{T}]^{T}$, and we have $\mathbf{h}_{RI,k}^{H}\bm{\Phi}\mathbf{x}_{k}=\sum_{g\in\mathcal{G}}\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{x}_{k,g}$. We then apply the vectorization property to transform $\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{v}_{p,g}$ and $\mathbf{h}_{RI,k,g}^{H}\bm{\Phi}_{g}\mathbf{x}_{k,g},~\forall g\in\mathcal{G}$, as
@@
The challenge of solving problem (60) is equivalent to that of solving the problem (19) in the SISO case, as both are subject to the lossy BD-RIS constraints (17b) – (17d). The core to deal with these constraints in both problems is the ADMM framework, where the same procedure described in Section III-B.2 is applied to solve (60). The only minor difference lies in the subproblem of variable $\bar{\bm{\phi}}$, which is still an unconstrained convex optimization and admits a closed-form solution given by
@@
IV-E Complexity Analysis
In this subsection, we provide a complexity analysis for the BCD iterative algorithm. In each iteration, the update of $\bm{\nu}$} and $\bm{\tau}$} require $\mathcal{O}\{K(M^{2}+MN+KN)\}$ operations, and the update of $\mathbf{W}$ has a complexity of $\mathcal{O}\{KN^{2}+I_{\mathrm{bs}}N^{3}\}$, where $I_{\mathrm{bs}}$ is the number of iterations for the bisection search. For the update of $\bm{\Phi}$, the calculation of $\mathbf{Q}$ incurs a complexity of $\mathcal{O}\{K(MN+N^{2})+\max(K^{2}\bar{M}^{4},K\bar{M}M)+\bar{M}^{3}M^{3}\}$, and each iteration of updating $\bar{\bm{\phi}}$, $\bar{\mathbf{y}}$, and $\mathbf{z}$ involves $\mathcal{O}\{\bar{M}^{3}M^{3}\}$ operations. Hence, the overall complexity of the BCD algorithm is $\mathcal{O}\{I(I_{\mathrm{bs}}N^{3}+KMN+\max(KM^{2},K^{2}\bar{M}^{4})+\max(K^{2}N,KN^{2})+I^{\prime}\bar{M}^{3}M^{3})\}$, where II and $I^{\prime}$ denote the number of iterations for BCD and updating $\bm{\Phi}$, respectively. The aforementioned complexity remains unchanged when adapted for forest-connected BD-RIS.
@@ V Performance Evaluation
In this section, we present the simulation results to evaluate the impact of losses in the group- and forest-connected reconfigurable admittance networks on the performance of BD-RIS-aided SISO and MU-MISO systems. The channels are modeled as an integration of large-scale fading and small-scale fading. Regarding the large-scale fading, we use the distance-dependent pathloss model given by $PL_{s}=\zeta_{0}(d_{s}/d_{0})^{-\varepsilon_{s}}$, $\forall s\in\{\mathrm{RT,RI,IT}\}$. Here, $\zeta_{0}$ denotes the path loss at the reference distance $d_{0}$, $d_{s}$ represents the distance associated with each link ($\mathrm{RT,RI,IT}$: transmitter-to-receiver, RIS-to-receiver, and transmitter-to-RIS, respectively), and $\varepsilon_{s}$ corresponds to the path loss exponent for each link. For small-scale fading, the transmitter-to-receiver link is modeled as a Rayleigh fading channel, whereas the RIS-to-receiver and transmitter-to-RIS links are modeled as Rician fading channels. The Rician factors for these links are denoted by $\kappa_{i}$, $\forall i\in\{\mathrm{RI,IT}\}$. In the following simulations for both SISO and MU-MISO systems, we set $\zeta_{0}=-30$ dB at a reference distance $d_{0}=1$ m for all channels. The distances are given as $d_{\mathrm{RT}}=52$ m, $d_{\mathrm{RI}}=2.5$ m and $d_{\mathrm{IT}}=50$ m. The path loss exponents are set to $\varepsilon_{\mathrm{RT}}=3.8$, $\varepsilon_{\mathrm{RI}}=2.2$ and $\varepsilon_{\mathrm{IT}}=2.5$. The Rician factors for the Rician fading channels are fixed at $\kappa_{\mathrm{RI}}=\kappa_{\mathrm{IT}}=2$ dB. The noise power is assumed as $\sigma^{2}=-80$ dBm for SISO systems, and $\sigma^{2}_{k}=-80$ dBm, $\forall k\in\mathcal{K}$, for MU-MISO systems.
(a) $M$-ADMM algorithm
@@
Figure 3: (a): Average rate versus $L_{1}$ for SISO systems ($R=2.5\ \Omega$, $M=32$, $\bar{M}=\{1,4,32\}$, $P=20$ dBm); (b): Sum-rate versus $L_{1}$ for MU-MISO systems ($R=3\ \Omega$, $N=K=4$, $M=32$, $\bar{M}=\{1,4,32\}$, $P=20$ dBm).
In Fig. 3, we first illustrate the selection of the value for $L_{1}$ by plotting the average rate and sum-rate performance against different practical $L_{1}$ values, using $M$-ADMM algorithm for SISO systems and BCD algorithm for MU-MISO systems, respectively. A practical range of parasitic inductance values is considered, and the performance of various BD-RIS architectures is evaluated within this range. For tree-connected and forest-connected architectures, the tridiagonal form is adopted. The results demonstrate that the value of $L_{1}$ in the circuit model has a significant impact on performance. Specifically, $L_{1}=6$ nH is identified as a reasonable choice, providing generally good performance in both SISO and MU-MISO systems. This value of $L_{1}$ is therefore used in all subsequent simulations [^11] The values of $L_{2}$ and $C_{m_{g},n_{g}},~\forall m_{g},n_{g}$, in the simulations are determined from the datasheet of the SMV2020-079LF (see Section II-B), while the value of $R$ is varied to investigate its impact on performance.. The performance differences among BD-RIS architectures will be discussed in the next two subsections (V-B and V-C). From an optimization perspective, $L_{1}$ influences the vertical shift of the circle defined by (9) for each tunable admittance component, thereby affecting their feasible admittance values during optimization. From a practical design perspective, a 66 nH inductor has sufficiently low inductance to avoid introducing non-negligible parasitic resistance in the circuit [Murata_LQG15WZ].
V-A Convergence Behavior of the Proposed Algorithms
In Figs. 46, we evaluate the convergence behavior of the proposed $M$-ADMM, low-complexity, and BCD algorithms, where all of these algorithms are terminated when the relative difference between two consecutive iterations is less than $10^{-4}$. As shown, all three algorithms exhibit stable convergence. In particular, the double-loop $M$-ADMM and BCD algorithms converge within fewer than 30 outer iterations. Although each outer iteration requires an inner loop of about 500 iterations to update $\bm{\Phi}$, the small number of outer iterations ensures that the overall algorithms remain computationally efficient. Compared to $M$-ADMM, the low-complexity algorithm involves only a single loop and converges more rapidly due to its much simpler objective function.
(a) Outer loop of the $M$-ADMM algorithm
(b) Inner loop of the $M$-ADMM algorithm
Figure 4: (a) Convergence of the proposed $M$-ADMM algorithm; (b) Convergence of the ADMM algorithm in the inner loop of $M$-ADMM algorithm ($R=1\ \Omega$, $M=30$, $\bar{M}=\{1,3,30\}$, $P=20$ dBm).
Figure 5: Convergence of the low-complexity algorithm ($R=1\ \Omega$, $M=30$, $\bar{M}=\{1,3,30\}$, $P=20$ dBm).
(a) Outer loop of the BCD algorithm
(b) Inner loop of the BCD algorithm
Figure 6: (a) Convergence of the BCD algorithm; (b) Convergence of the $\bm{\Phi}$-subproblem update within the BCD framework ($R=1\ \Omega$, $N=K=4$, $M=32$, $\bar{M}=\{1,8,32\}$, $P=20$ dBm).
V-B BD-RIS-aided SISO System
(a) $M$-ADMM algorithm